a boy pulls a rope attatched to a sled with a force of 80.0 N. The

rope makes an angle of 30 degrees with the ground. compute the effective value of the pull tending to move the sled. a) along the ground b)lifting it vertcally


1
Expert's answer
2021-05-25T10:10:11-0400

Force of pulling, "F=80\space N"

Angle with the horizontal, "\theta=30\degree"

Force along the ground = "F\cos\theta=80\cos30\degree=69.28\space N"

Force tending to lift it vertically = Force along y direction "=F\sin\theta"

"=80\sin30\degree=40\space N"